# Semi-Supervised Preference Optimization with Limited Feedback

### v2

@@ C.3 Real-Data Experiments
Table 8 describes a detailed configuration of experiments on real-world datasets. Our implementation builds upon the LlamaFactory framework (zheng2024llamafactory), a widely used open-source library for training, fine-tuning, and serving LLMs. We detail the hyperparameter configurations used for each baseline method across the three backbone models. All experiments were conducted using 2 GPUs of A6000, with tuning strategies adjusted per model size. All unspecified hyperparameters for Mistral and Llama3 baselines follow the configurations provided in SimPO [^2] https://github.com/princeton-nlp/SimPO (meng2024simposimplepreferenceoptimization).
@@
We use LoRA (hu2021loralowrankadaptationlarge) for rank 8 after exploring ranks from the set {4, 8, 32, 256} for all backbones. We perform a learning rate sweep between 1e-4 and 1e-9 and report results with the best performing value. The total batch size, gradient accumulation steps, and cutoff length are fixed at 64, 8, and 1024, respectively, across all models. We select the best model based on the minimum validation loss.
@@ Appendix D The Difference between SSRM, SPA, and SSPO
In this section, we provide a detailed comparison between SSPO and two prominent semi-supervised alignment frameworks: Semi-Supervised Reward Modeling (SSRM(he-etal-2024-semi-supervised) and Spread Preference Annotation (SPA(kim2025spreadpreferenceannotationdirect). While all three methods aim to leverage unlabeled data to mitigate the scarcity of human preference labels, SSPO distinguishes itself through its theoretical grounding, data efficiency, and non-iterative training paradigm. We analyze these methods across four key dimensions: Data Composition, Algorithmic Formulation, Theoretical Foundation, and Computational Efficiency.
@@ D.1 Data Composition & Annotation
  • SPA: Utilizes a small seed of paired data $D_{L}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n_{L}}$ and assumes that there exists a set of unseen prompts $X_{t}=\{x_{u}^{(t,i)}\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration. It explicitly generates multiple responses for the unseen prompts during training to form new unlabeled pairs $D_{U}^{(t)}=\{(x_{u}^{(t,i)},y_{u1}^{(t,i)},y_{u2}^{(t,i)})\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration, which are then self-labeled using the model’s implicit reward by DPO (rafailov2024direct).

### v3

@@ C.3 Real-Data Experiments
Table 8 describes a detailed configuration of experiments on real-world datasets. Our implementation builds upon the LlamaFactory framework (Zheng et al., 2024), a widely used open-source library for training, fine-tuning, and serving LLMs. We detail the hyperparameter configurations used for each baseline method across the three backbone models. All experiments were conducted using 2 GPUs of A6000, with tuning strategies adjusted per model size. All unspecified hyperparameters for Mistral and Llama3 baselines follow the configurations provided in SimPO [^2] https://github.com/princeton-nlp/SimPO (Meng et al., 2024).
@@
We use LoRA (Hu et al., 2021) for rank 8 after exploring ranks from the set {4, 8, 32, 256} for all backbones. We perform a learning rate sweep between 1e-4 and 1e-9 and report results with the best performing value. The total batch size, gradient accumulation steps, and cutoff length are fixed at 64, 8, and 1024, respectively, across all models. We select the best model based on the minimum validation loss.
@@ Appendix D The Difference between SSRM, SPA, and SSPO
In this section, we provide a detailed comparison between SSPO and two prominent semi-supervised alignment frameworks: Semi-Supervised Reward Modeling (SSRM(He et al., 2024) and Spread Preference Annotation (SPA(Kim et al., 2025). While all three methods aim to leverage unlabeled data to mitigate the scarcity of human preference labels, SSPO distinguishes itself through its theoretical grounding, data efficiency, and non-iterative training paradigm. We analyze these methods across four key dimensions: Data Composition, Algorithmic Formulation, Theoretical Foundation, and Computational Efficiency.
@@ D.1 Data Composition & Annotation
  • SPA: Utilizes a small seed of paired data $D_{L}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n_{L}}$ and assumes that there exists a set of unseen prompts $X_{t}=\{x_{u}^{(t,i)}\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration. It explicitly generates multiple responses for the unseen prompts during training to form new unlabeled pairs $D_{U}^{(t)}=\{(x_{u}^{(t,i)},y_{u1}^{(t,i)},y_{u2}^{(t,i)})\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration, which are then self-labeled using the model’s implicit reward by DPO (Rafailov et al., 2024).